OpenStudy (anonymous):

If XY+20=YW and XY=DT then DT+20=YW What property is this

OpenStudy (anonymous):

you are "substituting" using the knowledge that XY = DT to replace XY in the first equation with DT in the last equation... so I believe it is "Substitution"

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

Does that make sense? You have to be careful not to confuse it with "transitive"

OpenStudy (anonymous):

I honestly have no Idea on these types of questions...

OpenStudy (anonymous):

Let's replace those with the letters a, b, and c so the equations are easier to read: If a+20=b and a=c Then c+20=a

OpenStudy (anonymous):

Okay Thanks!

OpenStudy (anonymous):

Sorry, for the last equation, I meant to write: c+20=b